OpenStudy (anonymous):

i need help getting thru 8 system of equation poblems pls

OpenStudy (anonymous):

y=2x+1 -1/2y=x-1/2 the (- ) Signs are negatives!

OpenStudy (anonymous):

do you mean \[-\frac{1}{2}*y = x - \frac{1}{2}\] for the second equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what you should then is double the second equation to get \[-1 * y = 2x -1\] or \[-y = 2x -1\] which gives \[y = -2x +1\] you can then substitute this back into eqn 1 to give \[2x + 1 = -2x +1\] and so x = 0 (it is the only way -2x = 2x). then use this to sove for y, so you have y = 2(0) + 1, and so y = 1. then final solutions are x = 0, y = 1

OpenStudy (anonymous):

Y=3X+7 y=3X-6

OpenStudy (phi):

If you plot both of those lines, where do they intersect?